Please help!
One painter works 1 1/2 times as fast as another. Find their individual rates for painting a room if together it takes them 4 hours.

Answer by suresh(20)

Let x be the working rate of first painter and y be the rate of second painter.
Given:
x = (3/2) y
If they together work,they will complete the work in 4 hours.
For one hour first painter works 1/x and second painter 1/y
i.e. 1/x + 1/y = 1/4
2/3y+1/y=1/4

5/3y = 1/4
1/y=3/20
For second painter can do the work at the rate 20/3.
For first painter can do the work at the rate (3/2)*(20/3)=10